### v1

# Robust Hallucination Detection in LLMs via Adaptive Token Selection

Abstract
Hallucinations in large language models (LLMs) pose significant safety concerns that impede their broader deployment. Recent research in hallucination detection has demonstrated that LLMs’ internal representations contain truthfulness hints, which can be harnessed for detector training. However, the performance of these detectors is heavily dependent on the internal representations of predetermined tokens, fluctuating considerably when working on free-form generations with varying lengths and sparse distributions of hallucinated entities. To address this, we propose HaMI, a novel approach that enables robust detection of hallucinations through adaptive selection and learning of critical tokens that are most indicative of hallucinations. We achieve this robustness by an innovative formulation of the Hallucination detection task as Multiple Instance (HaMI) learning over token-level representations within a sequence, thereby facilitating a joint optimisation of token selection and hallucination detection on generation sequences of diverse forms. Comprehensive experimental results on four hallucination benchmarks show that HaMI significantly outperforms existing state-of-the-art approaches.
1 Introduction
Recent progress in Large Language Models (LLMs) has demonstrated impressive capabilities across a wide range of applications. However, the ever-growing popularity of LLMs also gives rise to concerns about the reliability of their outputs Ji et al. (2023); Chuang et al. (2023). Some research has indicated that LLMs are susceptible to hallucinations, which can be described as unfaithful or incorrect generations Longpre et al. (2021); Adlakha et al. (2023); Zhang et al. (2023). This tendency not only impedes the broader applications of LLMs but also poses potential safety risks, especially in high-stake fields such as legal and medical services. Therefore, the reliable detection of hallucinations is critical for the safe deployment of LLMs.
Various approaches have been developed to detect hallucinations. Recent studies indicate that predictive uncertainty can serve as useful detection features Kadavath et al. (2022); Burns et al. (2022); Duan et al. (2024), as predictions with low confidence often correlate with the presence of hallucinated content. Some research focuses on the evaluation of a single generation Manakul et al. (2023) while some harness semantic information from several samples, e.g., the semantic equivalences across multiple generations for the same question Farquhar et al. (2024). The latter is to exploit the observation that LLMs would consistently generate the same response if they were certain of it. Nevertheless, these methods are mainly based on the final generation output, rendering them ineffective in leveraging important semantics in the internal state representations.
Figure 1: Tokens that contain the most sufficient information related to correctness may appear at various positions within the sequence.
Another line of research focuses on the utilisation of the internal states of LLMs for hallucination detection. These internal representations can encode significant information about truthfulness direction of the generations. Motivated by this, great efforts have been made to explore the characteristics of these representations and train a binary classifier on them. Various supervision signals, including accuracy labels Li et al. (2024), converted semantic entropy labels Kossen et al. (2024), and eigenvalue-related labels Du et al. (2024), have been harnessed to train the classifier for detection tasks. One major challenge for these methods is that the majority of tokens in an incorrect/hallucinated response may not contribute to truthfulness. To address this issue, most of these methods focus on the use of predetermined tokens, such as the first generated token, the last generated token, or the one before the last. However, the exact location of the most indicative tokens for hallucination can vary significantly for generations of different questions, as illustrated in Figure 1, since the generation responses are of free-form with varying lengths and have sparse distributions of hallucinated entities. As a result, they can overlook important tokens where hallucinated information is actually concentrated.
To address this challenge, we propose HaMI, namely, Hallucination detection as Multiple Instance learning, an end-to-end joint token selection and hallucination detection approach that enables adaptive token selection from internal state representations for stable and accurate hallucination detection on generation responses of varying length. In HaMI we reformulate the task as a multiple instance learning (MIL) problem Carbonneau et al. (2018), where each response sequence is treated as a bag of token instances, with a bag-level label as either hallucinated (positive) or trustworthy (negative), and the objective becomes binary classification of the token bags. This way takes advantage of the fact that only a few token instances in the positive bag are positive since hallucinated content typically manifests in only a small subset of tokens within a sequence, whereas all token instances in the negative bag are always negative. In doing so, the MIL approach enables the exploitation of the hallucination labels at the sequence (bag) level to adaptively select the most responsive tokens for sequence-level hallucination detection.
To be more specific, LLMs are first prompted to generate response sequences with varying length. The MIL-driven hallucination detector is then optimised in HaMI to assign hallucination scores to all individual token instances and adaptively select the most indicative tokens in both positive and negative bags for the sequence-level prediction. The optimisation results in a detector that can distinguish the selected most positive hallucinated token instances from the hard negative instances (i.e., the tokens that have highest hallucination scores in a negative bag). Additionally, recognizing that predictive uncertainty serves as an important indicator of correctness, we further propose a representation enhancement module in HaMI, where we integrate multiple levels of uncertainty information into the original representation space for more effective training of our HaMI detector.
In summary, our contributions are as follows:
  • We propose a novel MIL-based framework HaMI for hallucination detection, which enables an end-to-end joint optimisation of token selection and hallucination detection. This results in adaptive hallucinated token selection, effectively mitigating the performance instability on response generations of varying length and hallucination entities. To our best knowledge, this is the first approach allowing such a joint optimisation.
  • We further introduce a module that incorporates internal representations with uncertainty scores to provide more hallucination indication information for the join optimisation in HaMI.
  • Comprehensive empirical results with widely adopted LLMs on four popular benchmark datasets show that HaMI can significantly outperform state-of-the-art (SOTA) competing methods.
Figure 2: The framework of our proposed HaMI. The LLM is prompted to generate answer tokens accompanied by token representations $h_{i}$. The network receives sequences of token representations from both positive $\mathcal{B^{+}}$ and negative $\mathcal{B^{-}}$ bags as inputs. The hallucination detector initially extracts high-level representations, followed by assigning a hallucination score to each token instance. We choose the largest scores from both positive and negative bags, subsequently maximizing the discriminative margin between them by minimizing a MIL loss as described in Eq. 4. Given the sequential nature of the generations, a constraint on the smoothness of hallucination scores of adjacent tokens is also added into HaMI.
2 Related Work
The term hallucination, under the closed-book setting, can refer to unfaithful or fabricated generations Zhang et al. (2023); Ji et al. (2023). Having gained wide interest, hallucination detection is crucial for LLMs to maintain high reliability in various specific tasks. These methods can be categorised as two main lines: uncertainty measurement and internal state analysis.
Uncertainty measurement.
Uncertainty measurement has been widely explored for hallucination detection. Some research focuses on token-level uncertainty Talman et al. (2023); Duan et al. (2024) with the assumption that low predictive logits or high entropy over tokens’ predictive distribution indicates the high possibility of hallucination. Some studies quantify sentence-level uncertainty by instructing LLMs to express the predictive uncertainty themselves with prompts like “Is your answer True or False?” Kadavath et al. (2022); Lin et al. (2022); Zhou et al. (2023); Manakul et al. (2023). Furthermore, there are many studies exploring exploiting semantic equivalence by measuring the consistency among multiple responses sampled from LLMs Mündler et al. (2023); Dhuliawala et al. (2023). For example, Farquhar et al. (2024) proposes Semantic Entropy, which employs a powerful LLM to evaluate semantic entailment among multiple generations and calculate semantic entropy over the entailment identity as uncertainty scores. Nonetheless, the performance of these methods relies on external tools and ignores the important semantics embedded in the internal representations of LLMs.
Internal state analysis.
Recently, a branch of works suggests that the internal states of LLMs encode more knowledge than they express and can reveal truthfulness direction Hubinger et al. (2024); Chen et al. (2024). The majority of this line employs probes Alain (2016) to better understand layer-wise representations and predict the correctness of generations Li et al. (2024); Marks and Tegmark (2024). Some research extends these methods by proposing new supervision signals. For example, Du et al. (2024) performs singular value decomposition on internal representations and calculates the norm of these representations projected on singular vectors as class scores, which are converted to class labels based on insights from a small set of wild data. Kossen et al. (2024) argue that the aforementioned semantic entropy value is preferable to accuracy labels for supervised training. While most of these works leverage predefined token representations, the truthfulness information is concentrated in specific tokens. Orgad et al. (2024) attempt to prompt an LLM to find the exact token in the sequence, but this method is resource-intensive and the reliability of detection is greatly impressed by the capability of the employed LLM. Instead of employing another LLM to find the unique exact answer, we train a ranking model to automatically select the critical tokens for hallucination detection in an end-to-end manner.
3 Preliminaries
Given a sequence of input tokens $\mathbf{x}=\{x_{1},x_{2},\dots,x_{n}\}$ consisting of a specific question $q$, an LLM will generate a sequence of tokens $\mathbf{y}=\{y_{1},y_{2},\dots,y_{t}\}$. Generally, each token $y_{i\in\{1,2,\dots,t\}}$ is decoded from the next predictive distribution over the model’s vocabulary set $\mathcal{V}$, formulated as $y_{i}=\arg\max_{y\in\mathcal{V}}P(y|y_{<i},\mathbf{x})$, and the predictive probability, i.e., logit, is denoted by $p_{i}$ for short. For each $\mathbf{y}$, apart from the token level uncertainty $p_{i}$, we can also obtain the overall uncertainty of the entire sequence, presented as
$p_{s}=(\prod_{i=1}^{t}p_{i})^{-\frac{1}{t}}.$ (1)
By accessing the internal states of the model, we can extract the internal representation $h_{i,l}\in\mathbbm{R}^{d}$ at layer $l$ for each token $y_{i}$, where $d$ is decided by the dimensions of the internal states of the LLM. The sequence feature space is denoted as $\mathcal{H}.$ The correctness of generated response is evaluated by GPT-4 Achiam et al. (2023) with label $z\in\{0,1\}$.
Given a datset $D=\{(q_{n},a_{n}\}_{n=1}^{N}$, where $\{q_{n}\}_{n=1}^{N}$ and $\{a_{n}\}_{n=1}^{N}$ are questions and ground-truth answers respectively, LLMs will generate answers $\mathbf{y}_{n}$ accompanied with predictive logits, hidden representations =$\mathcal{H}=\{\mathbf{h}_{n}\}_{n=1}^{N}$ and correctness labels $z_{n}\in\{0,1\}$ evaluated by GPT-4 Achiam et al. (2023). Our MIL-based hallucination detection is to automatically identify the most representative tokens in incorrect responses and the hard negative tokens (i.e., the most likely hallucinated tokens within correct responses) for training stable and accurate detectors.
4 Methodology
Our proposed HaMI aims to distinguish hallucination-free and hallucination-containing text generations by an adaptive token selection approach implemented by a joint optimisation of token selection and hallucination detection. Additionally, we introduce a predictive uncertainty-based module to integrate more hallucination features in HaMI and enhance its discriminative capability. The overall framework is presented in Figure 2. Below we introduce these two modules in detail.
4.1 MIL-driven Adaptive Token Selection
For a given sentence, its correctness is often determined by just a few words, such as noun entities Chuang et al. (2023), suggesting that truthfulness information may be encoded in the internal representations of specific tokens. Nevertheless, the correctness label is typically assigned to the entire sequence. The key idea of our approach is to adaptively identify the most salient tokens, upon which we can train a reliable hallucination detector. To achieve this goal, we introduce Multiple Instance Learning (MIL) to this domain.
In MIL, instead of finding the classification boundary for samples with different identities, it tries to distinguish between the hard instance from sample bags of various categories. To make it clear, there are positive bags and negative bags containing multiple instances. All the instances in negative bags are negative while only a few instances in positive bags are positive. MIL aims to separate the positive instances and the hard negative instances from negative bags. This objective aligns with our assumption that there are only several tokens containing information of hallucination.
Therefore, we reformulate hallucination detection with adaptive token selection as a MIL problem. Particularly, the generated sequence can be regarded as a bag of tokens. Generations without hallucinations are labelled as negative bags $\mathcal{B^{-}}$ (label ‘0’), while those with hallucinations are labelled as positive bags $\mathcal{B^{+}}$ (label ‘1’). The representations of the positive and negative sequences are denoted as $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$ respectively and $\mathbf{h}=\{h_{i}\}_{i=1}^{t}$. Intuitively, the detector is expected to assign higher scores to token instances from the positive bag compared to those from the negative bag. Given the hallucination sparsity insight mentioned before, we choose the token instance with the highest score as the salient token, which can be defined as
$i^{+}=\arg\max\limits_{i}f_{\theta}(\mathbf{h}^{+}),$ (2)
$i^{-}=\arg\max\limits_{i}f_{\theta}(\mathbf{h}^{-}),$ (3)
where $i\in\{1,2,\cdots,t\}$, $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$ are drawn from input representation space $\mathcal{H}$ w.r.t. $\mathcal{B^{-}}$ and $\mathcal{B^{-}}$ respectively. As such, we are able to locate the instance most likely to represent a true positive in the positive bag and the most challenging instance resembling a hallucination in the negative bag as aforementioned salient tokens. Our approach seeks to distinguish between positive and negative samples by maximizing the distance between the selected instances from these two categories in representation space. The training MIL objective can then be formulated in a hinge-loss style as follows:
$\mathcal{L}_{MIL}=max(0,1-\max_{n\in\mathcal{B^{+}}}f_{\theta}(\mathbf{h}_{n}^{+})+\max_{n\in\mathcal{B^{-}}}f_{\theta}(\mathbf{h}_{n}^{-}),$ (4)
where $\theta$ are parameters of the hallucination detector as presented in Figure 2. In doing so, the adaptive token selection can mitigate the limitations caused by predefined token location. For example, the commonly used First generated token is hard to capture hallucination appears in the subsequent generations, especially when the generation is long or LLMs react to the input politely.
Note that the next token is conditioned on all previous tokens, so the token generation process contains a sequential nature. As depicted in Figure 2, on the same side of the peak, the hallucination scores for two adjacent tokens tend to be more similar. Therefore, we exploit sequential smoothness via the following loss
$\mathcal{L}_{Smooth}=(f_{\theta}(h_{i})-f_{\theta}(h_{i-1}))^{2},$ (5)
through which we aim to ensure the consistency of the hallucination scores of neighbouring tokens. Therefore, the adaptive token selection is achieved by minimizing the following overall loss:
$\mathcal{L}_{ATS}=\mathcal{L}_{MIL}+\mathcal{L}_{Smooth}.$ (6)
4.2 Augmenting Internal State Representations with Predictive Uncertainty
A lot of studies have demonstrated that uncertainty measurements are effective for hallucination detection. Our approach seeks to determine whether incorporating uncertainty metrics into the original token representation space can augment these representations’ ability to distinguish truthfulness.
The collection of predictive uncertainty measurements can be categorised into three levels: i) token-level uncertainty, i.e., predictive logits:
$p_{i}=P(y|y_{<i},\mathbf{x}),$ (7)
where $\mathbf{x}$ are prompt tokens and $y$ are generated tokens; ii) sentence-level perplexity, which is monotonically related to the mean of the negative log-likelihood of the sequence:
$P_{s}=-\frac{1}{T}\sum_{t=1}^{T}\log P(y|y_{<i},\mathbf{x}),$ (8)
and iii) semantic consistency across multiple samples, where the uncertainty value can be quantified by the number of semantic-equivalence generations over the whole generations based on the entailment results Farquhar et al. (2024). Specifically,
$P_{c_{m}}=\sum_{1}^{M}\frac{I_{c}=c_{m}}{M},$ (9)
where $M$ is the total number of generations, $c_{m}$ is the identified cluster and $I_{c}$ is the assigned cluster identity of a given sample. These uncertainty metrics can be directly leveraged for hallucination detection, and on average, the semantic consistency outperforms the other two metrics. However, it requires more computational cost while the other two metrics can obtained from just one generation as long as we can access the internal states of LLMs.
To augment the internal representations with the uncertainty information, we define the final input representation for each token as follows:
$\mathbf{h}^{\prime}=(\lambda_{1}+\lambda_{2}\cdot P_{uncertainty})\cdot\mathbf{h},$ (10)
where $\lambda_{1}$ and $\lambda_{2}$ are used to control the impact of uncertainty metrics. The improvements gained by various uncertainty measurements are evaluated and discussed in Section 5.4.
LLaMA-2-chat-7B LLaMA-2-chat-13B
Trivia QA SQuAD NQ BioASQ Trivia QA SQuAD NQ BioASQ
SE Farquhar et al. (2024) 0.879 0.799 0.801 0.823 0.777 0.767 0.769 0.743
p(True) Kadavath et al. (2022) 0.644 0.609 0.533 0.569 0.580 0.579 0.565 0.612
Perplexity Ren et al. (2023) 0.747 0.634 0.683 0.594 0.758 0.683 0.691 0.707
RMD Ren et al. (2023) 0.531 0.525 0.555 0.603 0.530 0.543 0.519 0.566
HaloScope Du et al. (2024) 0.625 0.574 0.615 0.682 0.564 0.543 0.575 0.624
LP-First Li et al. (2024) 0.796 0.760 0.715 0.823 0.690 0.699 0.670 0.720
LP-Last Kossen et al. (2024) 0.826 0.755 0.741 0.739 0.689 0.724 0.745 0.639
HaMI (Ours) 0.923 0.812 0.823 0.845 0.839 0.783 0.778 0.792
Table 1: AUCROC results of HaMI and seven SOTA competing hallucination detectors on four datasets with two LLaMA-based LLMs of different size. The results with Mistral are shown in the supplementary due to space limitation.
5 Experiments
We evaluate HaMI across various datasets and models. In this section, we first illustrate the experimental setup in Section 5.1 and present the main results in Section 5.2. We also conduct in-depth studies to analyse the robustness of our approach in Section 5.3 and the effectiveness of various components in Section 5.4.
5.1 Setup
Datasets and models.
We investigate our methods on four question-answering (QA) datasets across a range of domains, including (1) Trivia QA, a relatively complicated confabulation QA datasets revealed by Joshi et al. (2017); (2) Stanford Question Answering Dataset (SQuAD for short) Rajpurkar (2016), which is based on Wikipedia and generated by humans through crowdsourcing; (3) Natural Questions (denoted as NQ) Kwiatkowski et al. (2019), containing search information from real users on Google search and (4) a biomedical QA corpus BioASQ Krithara et al. (2023). For each dataset, we randomly extract $3,000$ QA pairs for training and $800$ pairs for testing. Consistent with multi-sampling approaches, we prompt LLMs six times for each question for testing data. Following Farquhar et al. (2024), we have $400$ questions randomly sampled from the generated test set at the testing stage for evaluation.
[…]
Baselines.
We compare our method with a series of baselines, covering uncertainty-based methods and internal representation-based methods. The uncertainty-based methods are as follows: Semantic Entropy (SE), semantic-equivalence measurement across multiple samples Farquhar et al. (2024), which employ GPT-3.5 for consistency evaluation; p(True) Kadavath et al. (2022), asking LLMs to express the probability or correctness for given answers themselves; Perplexity Ren et al. (2023), which is based on the Eq. 8. For internal representation-based methods, we choose the method based on relative Mahalanobis distance (RMD) Ren et al. (2023) and probing classifiers with different settings on supervising signals and token locations. We employ HaloScope Du et al. (2024), which proposes an automated membership estimation score and converts the score to binary labels for classification. For various location selection strategies, we choose the first and the last token as settled in many works, denoted as LP-First and LP-Last respectively Li et al. (2024); Orgad et al. (2024); Azaria and Mitchell (2023); Marks and Tegmark (2024); Kossen et al. (2024).
Evaluation.
Following Farquhar et al. (2024); Ren et al. (2023), we evaluate the model’s capability for hallucination detection by calculating the area under the receiver operating characteristic curve (AUROC), which is a widely used metric to evaluate the discriminative ability of binary classification. The ground-truth labels indicating the correctness of answers are given by GPT-4 Achiam et al. (2023), which is prompted to determine if the answer is correct or not based on the consistency between generated answers and gold answers and their own knowledge. The prompt template refers to Farquhar et al. (2024). Note that since GPT-4 makes mistakes mostly for the positive samples as checked with the gold answers, we ask GPT-4 to rejudge samples labelled as positive and discard samples if the result is inconsistent with the first result.
5.2 Main Results
In Table 1, we evaluate our proposed HaMI by comparing it with seven competing SOTA detection methods on four diverse QA datasets in LLMs of two different sizes. As depicted in the table, our approach achieves superior performance compared with other methods in all cases. In particular, HaMI outperforms others by a large margin on the Trivia QA dataset in both LLaMA-2-chat-7B and LLaMA-2-chat-13B models. Notably, both p(True) and Semantic Entropy resort to external LLMs for assistance, but their capability for hallucination detection differs significantly. Following the configurations described in their respective original studies, p(True) employs the LLaMA-7B model, while Semantic Entropy harnesses the more powerful GPT-3.5 model. This variation underscores that the performance of uncertainty-based methods, which rely on external support, is highly dependent upon the capabilities of the selected assistant LLM. Furthermore, among internal representation-based methods, supervised approaches generally outperform the unsupervised ones. For example, LP-First, LP-Last and our approach exhibit superior performance than HaloScope. LP-First can achieve impressive performance on some datasets like BioASQ but fail to work well on the other datasets. On the other hand, LP-Last works well on datasets like Trivia QA and less effectively on the other datasets. These results showcases the unstable performance of predetermined token-based methods. By contrast, our proposed HaMI performs consistently well on all four benchmarks, surpassing all competing methods. This superiority is due to not only the predictive uncertainty enhancement but also the effectiveness of adaptive token selection on generation responses of diverse forms (see Section 5.4 for more detailed analysis).
Figure 3: AUROC results of cross-dataset generalisation on four datasets. Experiments are based on LLaMA-2-chat-7B.
5.3 Cross-dataset Generalisation Ability
The ability to generalise across datasets is essential to facilitating real-world applications of LLMs in diverse domains. We conduct experiments on the aforementioned four datasets to assess whether the proposed HaMI can effectively generalise among various datasets. For each dataset, we report the average AUROC scores of detectors trained on one of the other three datasets.
The cross-dataset generalisation performance is illustrated in Figure 3. It is clear that our method HaMI achieves consistently the best generalisation performance across all four datasets, outperforming the best competing method LP-First by 8% - 18%. Compared to the within-dataset performance in Table 1, the maximum performance decline in HaMI is no more than $4.0\%$, observed on the Trivia QA dataset, which is significantly lower than the competing methods. Notably, when trained on the other datasets, HaMI achieves an average AUROC score of $0.819$ on the SQuAD dataset, which marginally exceeds the score of $0.813$ when directly leveraging the SQuAD dataset as the training set. These findings affirm the effectiveness of HaMI as a reliable hallucination detector, demonstrating its adaptability in varied contexts.
5.4 Ablation Study
In this section, we comprehensively evaluate the effectiveness of different modules of our proposed HaMI.
Effectiveness of different modules.
In Table 2, we offer a detailed comparison of the improvements provided by various modules. The Baseline model is the one that uses the original internal representation of the First token as model inputs and does not use any of our modules. For adaptive token selection (ATS), we further explore the importance of the two loss functions, including both the MIL loss and the smoothness loss. Our results indicate that both of our ATS module and uncertainty-based semantic augmentation respectively contribute significant significant improvement over the baseline. In comparison to the augmentation module, the improvement from the ATS module is much larger. Within ATS, the smoothness loss can result in additional detection enhancement. Moreover, apart from the BioASQ dataset, substantial improvements are observed with the implementation of ATS following the integration of the semantic consistency score. This finding is especially crucial for practical applications, as the ATS module does not introduce additional computational burdens.
$P_{c_{m}}$ ATS TQ SQ NQ BQ
$\mathcal{L}_{MIL}$ $\mathcal{L}_{Smooth}$
0.799 0.714 0.656 0.780
0.858 0.756 0.724 0.828
0.858 0.785 0.795 0.803
0.906 0.787 0.776 0.839
0.923 0.812 0.823 0.845
Table 2: Experimental results on the effectiveness of the proposed uncertainty-augmented internal representations and the ATP module. Trivia QA, SQuAD, and BioASQ are denoted as TQ, SQ, and BQ for presentation respectively. Experiments are based on LLaMA-2-chat-7B.
Location Trivia QA SQuAD NQ BioASQ
First 0.858 0.756 0.724 0.828
Last 0.874 0.768 0.788 0.784
Mean 0.900 0.806 0.805 0.836
Ours 0.923 0.812 0.823 0.845
Table 3: AUROC results on the analysis of our adaptive token selection module. Experiments are based on LLaMA-2-chat-7B.
Figure 4: Adaptive token selection results showing the top two tokens with the largest hallucination scores.
In-depth analysis of adaptive token selection.
Given that hallucinations may appear occasionally in long-form generations, identifying the critical token is important for capturing sufficient truthful information. Our proposed adaptive token selection (ATS) module solves the above concern. We conduct experiments to investigate the performance of the proposed ATS module, comparing it against commonly used benchmarks such as the First generated token, the Last generated token, and the Mean of all tokens. The results in Table 3 demonstrate that our selection strategy outperforms the alternatives across all datasets, yielding an average improvement of $8\%$ over the First token and $4\%$ over the Last token respectively. Compared with the single selection strategy, the Mean token can achieve more significant performance. We notice that the First token achieves better performance on the BioASQ dataset while showing a compromised behaviour on the other three datasets. The Last token is just the opposite. Considering that the average number of generated tokens for each question of the BioASQ dataset (i.e., $56$) is larger than the left datasets (Trivia QA - $22$, SQuAD - $30$, NQ - $30$), it is concluded that for longer generation, the last token may fail to retain important truthful information captured by earlier parts of the sequence. Conversely, in shorter generations, the last token can capture full semantics while the first token may miss content in the subsequent predictions. This suggests that i) the performance of the previous predefined token location-based approach could be sensitive to the length of the text generated and ii) our ATS module can well mitigate this issue.
Figure 4 presents an illustrative example of the scoring results over tokens in a positive bag and a negative bag, where we can observe that there is significant distinguishability between the positive and negative tokens since the maximum scores of the instances in the positive bag is substantially greater than that in the negative bag. Tokens denoted by blue characters have scores that are very close to the largest ones and we find that these tokens can be adaptively identified by our method as concentrated answers in comparison to ground-truth answers. Additionally, it is noted that while selected tokens are associated with the exact answer, they may appear at any location in its vicinity, which can be captured by our smoothness loss.
Metrics Trivia QA SQuAD NQ BioASQ
Original $h_{i}^{l}$ 0.858 0.785 0.795 0.803
$\;\;+p_{i}$ 0.867 0.802 0.800 0.823
$\;\;+P_{s}$ 0.883 0.798 0.805 0.825
$\;\;+P_{c_{m}}$ 0.923 0.812 0.823 0.845
Table 4: AUROC results on the effectiveness of different uncertainty quantified metrics. $p_{i}$, $P_{s}$, and $P_{c_{m}}$ refer to logits, perplexity and semantic consistency respectively. Experiments are based on LLaMA-2-chat-7B.
Analysis of different uncertainty enhancement methods.
As shown in Table 2, integrating uncertainty metrics into the original representation space can enhance the distinguishability between correct and hallucinated generations. There are different methods for this integration, as discussed in Section 4.2. Here we systematically examine using different uncertainty measuring methods for the internal representation enhancement, including token-level logit $p_{i}$, sequence-level perplexity $P_{s}$, and semantic consistency $P_{c_{m}}$ across multiple samples, as detailed in Eq. 7, 8 and 9 respectively. The final inputs are derived using Eq. 10. Experimental results are reported in Table 4, where we can observe that all three uncertainty metrics contribute to improved detection capabilities over the baseline representations. Specifically, the enhancements attributed to semantic consistency across multiple generations are particularly significant, exhibiting improvements up to $8.3\%$. Although the improvements observed with the other two metrics are less pronounced, both methods surpass the performance of the commonly used binary classifier using original representations. Moreover, we notice that incorporating sequence-level uncertainty enables our method to achieve performance comparable to the SOTA multi-sampling approach, namely Semantic Entropy (SE), without incurring the costs associated with multiple generations and external LLM employments. This observation highlights the potential of HaMI for deployments in various practical environments that involve external tools or not.
Figure 5: Impact of different layers on hallucination detection. The results are based on LLaMA-2-chat-7B. The suffix ‘-ori’ denotes experimental results with original representations.
How does HaMI perform with representations from different layers?
We evaluate how layers impact the performance of HaML for hallucination detection with representations extracted from all 32 layers of the LLaMA-2-chat-7B model. Our investigation is based on the Trivia QA and NQ datasets, applying both original internal representations and semantic-equivalence-enhanced representations. As depicted in Figure 5, the AUROC values for using the original representations (indicated by the dashed line) exhibit a clear increase, peaking between layers 12 and 18, before declining to a relatively stable level between $0.75$ to $0.80$. This observation suggests that the truthfulness content evolves across the initial to middle layers. The performance of the uncertainty-enhanced representations, illustrated by solid lines, maintains a relatively consistent trend, with the highest AUROC scores concentrated in the middle layers. The comparison of these trends indicates that incorporating predictive uncertainty enhances the distinctiveness of the representations, particularly in the earlier layers where less semantic information is typically available.
6 Conclusion
Hallucination detection is essential for the reliable deployment of LLMs. In this paper, We introduce a joint token selection and hallucination detection approach, HaMI, designed to adaptively identify the most likely to be hallucinated tokens, enhancing the robustness of the detection on generation responses of varying lengths and hallucinated entities. Specifically, our approach incorporates a straightforward yet effective multiple instance learning formulation to automatically highlight salient tokens for the training of more accurate hallucination detectors. Additionally, we also explore integrating uncertainty metrics into the original representations to enrich them with more information about truthfulness. Extensive empirical results demonstrate that HaMI substantially outperforms existing methods across diverse QA datasets and LLMs with various characteristics. Our ablation studies offer additional investigations and insights into different aspects of designed modules. While our experiments primarily focus on the QA domain, the principles underlying our method are task-free, suggesting potential applicability to a broad spectrum of other tasks.